# The General Position Problem in Graph Theory: A Survey

### v3

@@ 2.3.2 Maximal outerplanar graphs
Outerplane graphs are planar graphs with a plane embedding such that every vertex lies on the boundary of the exterior region. An outerplane graph is maximal outerplane if adding any edge results in a non-outerplane graph. A straight linear 222-tree on $n$ vertices is a graph $G_{n}$ with vertex set $V(G_{n})=[n]$, where vertices $i$ and $j$ are adjacent in $G_{n}$ if and only if $0<|i-j|\leq 2$.
The paper [119] focuses on the general position number of maximal outerplane graphs and presents several key results. It is shown in [119, Lemma 2.8] that the general position number of a maximal outerplane graph is bounded in terms of the maximum degree as: $\operatorname{gp}(G)\geq\lfloor\frac{2(\Delta(G)+1)}{3}\rfloor$. In the next result, maximal outerplane graphs with $\operatorname{gp}(G)=3$ are characterised.
@@
[119, Theorem 3.1] If $G$ is a maximal outerplane graph with $n(G)\geq 7$, then $\operatorname{gp}(G)=3$ if and only if $G$ is a straight linear 222-tree.
In [119, Theorem 3.4], it was proved that for any maximal outerplane graph $G$ of order $n\geq 6$, $\operatorname{gp}(G)\leq\left\lfloor\frac{2n}{3}\right\rfloor$. The paper [119] then focuses on all maximal outerplane graphs that attain this bound.
@@ 2.3.3 Additional exact results
The general position numbers of glued binary trees $GT(r)$, $r\geq 2$, and of Beneš networks $BN(r)$, $r\geq 1$, were determined in [80] and [82] respectively. We do not give the definitions of these networks here, but recall that if $r\geq 2$, then $\operatorname{gp}(GT(r))=2^{r}$ [80, Proposition 3.8], and that if $r\geq 1$, then $\operatorname{gp}(BN(r))=2^{r+1}$ [82, Theorem 4.1]. Similar research has been done in [96, 97, 98]. In [96], the general position number is determined for hyper trees and shuffle hyper trees. In [97], the general position number is obtained for hexagonal networks, honeycomb networks, silicate networks, and oxide networks, while [98] reports the general position number of butterfly networks.
@@
[126, Theorem 2.5] If $G$ is a cactus graph with $k$ cycles and $t$ pendant edges, then $\operatorname{gp}(G)\leq\max\{3,2k+t\}$.
Moreover, [126, Theorem 2.5] also characterises the class of cactus graphs that attain equality in Theorem 2.18 as the cactus graphs in which all cycles are ‘good’; see the paper for the definition of a good cycle.
If $p\geq 3$ and $n\geq 1$, then the Sierpiński graph $S_{p}^{n}$ has vertex set $V(S_{p}^{n})=\{0,1,\ldots,p-1\}^{n}$, whilst vertices $i_{1}\cdots i_{n}$ and $j_{1}\cdots j_{n}$ are adjacent if there exists an index $h\in[n]$, such that (i) if $t<h$, then $i_{t}=j_{t}$, (ii) $i_{h}\neq j_{h}$, and (iii) if $t>h$, then $i_{t}=j_{h}$ and $j_{t}=i_{h}$, see [60]. In [104, Corollary 3.2] it is proved that if $p\geq 3$ then,
$\operatorname{gp}(S_{p}^{2})=\begin{cases}\frac{(p+1)^{2}}{4};&p\text{ odd}\,,\\[5.0pt] \frac{p(p+2)}{4};&p\text{ even}\,,\end{cases}\quad\text{and}\quad\#\operatorname{gp}(S_{p}^{2})=\begin{cases}\binom{p}{\frac{p+1}{2}};&p\text{ odd}\,,\\[5.0pt] \binom{p+1}{\frac{p+2}{2}};&p\text{ even}\,,\end{cases}$
where #$\#\operatorname{gp}(G)$ indicates the number of $\operatorname{gp}$-sets of a graph $G$. Moreover, [104, Corollary 4.5] asserts that if $n\geq 2$, then $\operatorname{gp}(S_{3}^{n})=3^{n-2}+3$.
@@ 2.4.1 Cartesian product
General position sets have been extensively researched so far on Cartesian products. As a first result, it was proved in [82, Corollary 3.2] that $\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4$, where $P_{\infty}$ is the two-way infinite path. In the same paper it was further proved that $10\leq\operatorname{gp}(P_{\infty}^{\operatorname{\,\square\,},3})\leq 16$, where $G^{\operatorname{\,\square\,},n}$ denotes the $n$-fold Cartesian product of $G$. These results were completed by the following result, where the Erdős-Szekeres Theorem on monotone sequences was applied to prove the upper bound.
@@
[68, Theorem 1] If $n\in{\mathbb{N}}$, then $\operatorname{gp}(P_{\infty}^{\operatorname{\,\square\,},n})=2^{2^{n-1}}$.
For general Cartesian products, we have the following lower bound proved in [51, Theorem 3.1], while the equality conditions are from [117, Theorems 3.7, 3.8]. A graph is a generalised complete graph if it is obtained by the join of an isolated vertex with a disjoint union of one or more complete graphs.
@@
If $G$ and $H$ are connected graphs, then
@@
The equality holds only if $\operatorname{gp}(G)\in\{n(G),n(G)-1\}$ and $\operatorname{gp}(H)\in\{n(H),n(H)-1\}$. Moreover, if $n(G)\geq 3$ and $n(H)\geq 3$, then the equality holds if and only if both $G$ and $H$ are generalised complete graphs.
@@
[117, Theorem 3.5] If $G$ is a connected, bipartite graph with $n(G)\geq 3$, then $\operatorname{gp}(G\operatorname{\,\square\,}K_{2})=\operatorname{gp}(G)$ if and only if G is a complete bipartite graph.
@@
[51, Theorem 3.2] If $k\geq 2$ and $n_{1},\ldots,n_{k}\geq 2$, then
@@
    [120, Theorem 2.1] $\operatorname{gp}(T\operatorname{\,\square\,}T^{\prime})=\operatorname{gp}(T)+\operatorname{gp}(T^{\prime})$.
@@
    [117, Theorem 3.2] $\operatorname{gp}(T\operatorname{\,\square\,}K_{m})=\operatorname{gp}(T)+\operatorname{gp}(K_{m})-1$.
From Theorem 2.21 we can deduce that if $r\geq 3$ and $s\geq 3$, then $\operatorname{gp}(P_{r}\operatorname{\,\square\,}P_{s})=4$. This result has been expanded as follows, where #$\#\operatorname{gp}(G)$ indicates the number of $\operatorname{gp}$-sets of a graph $G$.
@@
[66, Theorem 2.1] If $2\leq r\leq s$, then
#$\#\operatorname{gp}(P_{r}\operatorname{\,\square\,}P_{s})=\left\{\begin{array}[]{ll}\vskip 5.69054pt6;&r=s=2\,,\\ \vskip 5.69054pt\displaystyle{\frac{s(s-1)(s-2)}{3}};&r=2,s\geq 3\,,\\ \displaystyle{\frac{rs(r-1)(r-2)(s-1)(s-2)(r(s-3)-s+7)}{144}};&r,s\geq 3\,.\end{array}\right.$
@@
[73, Theorem 3.4] If $r$ and $s$ are integers with $r\geq s\geq 4$, then
@@
A formula for the general position number of the Cartesian product of a complete multipartite graph with a path is also given in [117, Theorem 3.6].
The $n$-dimensional hypercube $Q_{n}$, $n\geq 1$, is defined to be $K_{2}^{\operatorname{\,\square\,},n}$, the Cartesian product of $n$ copies of $K_{2}$. To determine $\operatorname{gp}(Q_{n})$ is a notoriously difficult problem. The only exact values known so far are $\operatorname{gp}(Q_{1})=\operatorname{gp}(Q_{2})=2$, $\operatorname{gp}(Q_{3})=4$, $\operatorname{gp}(Q_{4})=5$, $\operatorname{gp}(Q_{5})=6$, $\operatorname{gp}(Q_{6})=8$, and $\operatorname{gp}(Q_{7})=9$, see [73, Table 1]. Using a probabilistic construction, Körner [72] obtained general position sets in $Q_{n}$ of size $\frac{1}{2}\frac{2^{n}}{\sqrt{3^{n}}}$. He also noted that the problem of finding $\operatorname{gp}(Q_{n})$ is equivalent to finding the largest size of a (2,1)(2,1)( 2 , 1 )-separating system in coding theory. Defining
@@
[66, Theorem 5.1] If $G$ is a graph, then
@@ 2.4.2 Strong product and direct product
[71, Theorem 4.2, Corollary 4.1] If $G$ and $H$ are connected graphs, then
@@
[71, Proposition 4.3] If $G$ is a connected graph and $n\geq 1$, then $\operatorname{gp}(G\boxtimes K_{n})=n\cdot\operatorname{gp}(G)$. Moreover, if $\operatorname{gp}(G)=\omega(G\operatorname{{}_{SR}})$, then $\operatorname{gp}(G\boxtimes K_{n})=\omega((G\boxtimes K_{n})\operatorname{{}_{SR}})$.
@@ 2.4.3 Generalised lexicographic product
[71, Theorem 5.1] Let $G$ be a graph with $V(G)=\{g_{1},\ldots,g_{n}\}$ and let $k_{i}$, $i\in[n]$, be positive integers. If $S$ is a gp-set of $G$ that induces a complete subgraph of $G\operatorname{{}_{SR}}$, and $\min\{k_{i}:\ g_{i}\in S\}\geq\max\{k_{i}:\ g_{i}\notin S\}$, then
@@ 2.4.4 Sierpiński product
In [115], the Sierpiński general position number ${\rm gp}{{}_{\rm S}}(G,H)$, and the lower Sierpiński general position number $\underline{{\rm gp}}{{}_{\rm S}}(G,H)$, were introduced as the cardinality of a largest, resp. smallest, $\operatorname{gp}$-set in $G\otimes_{f}H$ over all possible functions $f\colon V(G)\rightarrow V(H)$, that is,
@@
[115, Theorem 3.2] If $G$ is a connected graph of order at least two and $u\in V(G)$, then
@@
[115, Theorem 4.2] If $G$ and $H$ are two connected graphs of order at least two, then
@@
If $G=K_{2}$ and $H$ is a connected graph with $n(H)\geq 2$, then $\underline{{\rm gp}}{{}_{\rm S}}(K_{2},H)=2\xi^{-}(H)$ and ${\rm gp}{{}_{\rm S}}(K_{2},H)=2\xi(H)$, see [115, Theorem 4.4]. Finally, for Sierpiński products of complete graphs the following is true.
@@ 2.5.1 Vertex- and edge-deleted subgraphs
[38, Theorem 3.1] If $x$ is a vertex of a graph $G$, then $\operatorname{gp}(G-x)\leq 2\operatorname{gp}(G)$. Moreover, the bound is sharp.
The sharpness examples for Theorem 2.38 provided in [38] are such that $G-x$ is not connected. Constructions are also given which demonstrate that $\operatorname{gp}(G-x)$ can be much larger than $\operatorname{gp}(G)$ even when $G-x$ is connected.
On the other hand, $\operatorname{gp}(G-x)$ cannot be bounded from below by a function of $\operatorname{gp}(G)$. For instance, consider the fan graphs $F_{n}$, $n\geq 3$, that is, the join between $P_{n}$ and $K_{1}$. Then $\operatorname{gp}(F_{n})=\lceil\frac{2n}{3}\rceil$ and $\operatorname{gp}(F_{n}-x)=2$, where $x$ is the vertex of $F_{n}$ of degree $n$. On the positive side, if $x$ lies in some $\operatorname{gp}$-set, then $\operatorname{gp}(G)-1\leq\operatorname{gp}(G-x)$, see[38, Proposition 3.3]. For edge-deleted subgraphs we have:
@@ 2.5.2 Complementary prisms
[87, Theorem 3.1] If $G$ is a connected graph, then $\operatorname{gp}(G\overline{G})\leq n(G)+1$, and if $G$ is disconnected, then $\operatorname{gp}(G\overline{G})\leq n(G)$.
In the next result graphs are characterised for which the first upper bound of Theorem 2.40 is achieved. To state it, we need the following two concepts, where the second represents the original idea that led to the $d$-positions sets which will be considered in Section 3.1. A vertex $v$ of a connected graph $G$ is a central vertex of $G$ if ${\rm ecc}_{G}(v)=\operatorname{rad}(G)$. A set $S$ of vertices in a graph $G$ is a 333-general position set if no three vertices from $S$ lie on a common shortest path of length at most three. By $N_{G}^{2}(v)$ we denote the set of vertices at distance two from $v$ in $G$.
@@
[87, Theorem 3.4] Let $G$ be a graph with $n(G)\geq 2$ and such that both $G$ and $\overline{G}$ are connected. Then $\operatorname{gp}(G\overline{G})=n(G)+1$ if and only if $\operatorname{rad}(G)=2$ and there exists a central vertex $v$ such that
@@
    $\overline{N_{G}(v)}$ is a 333-general position set in $\overline{G}$ and $N_{G}^{2}(v)$ is a 333-general position set in $G$, and
@@
[87, Theorem 4.1] If $G$ is a graph, then
@@
It follows from Theorem 2.42 that $\operatorname{gp}(G\overline{G})\geq\max\{\operatorname{gp}_{3}(G),\operatorname{gp}_{3}(\overline{G})\}$. In [87, Theorem 4.2] it was proved that among connected graphs the equality holds here if and only if $G$ is a complete multipartite graph. The paper [87] considers several additional classes of graphs $G$ and the corresponding values of $\operatorname{gp}(G\overline{G})$: bipartite graphs (in particular trees, hypercubes and Cartesian products of paths), split graphs and block graphs. For instance, if $T$ is a tree, then
$\operatorname{gp}(T\overline{T})=\begin{cases}n(G)+1;&\operatorname{diam}(T)=4\,,\\ n(G);&otherwise\,,\end{cases}$

### v4

@@ 2.3.2 Maximal outerplanar graphs
Outerplane graphs are planar graphs with a plane embedding such that every vertex lies on the boundary of the exterior region. An outerplane graph is maximal outerplane if adding any edge results in a non-outerplane graph. A straight linear $2$-tree on $n$ vertices is a graph $G_{n}$ with vertex set $V(G_{n})=[n]$, where vertices $i$ and $j$ are adjacent in $G_{n}$ if and only if $0<|i-j|\leq 2$.
The paper [136] focuses on the general position number of maximal outerplane graphs and presents several key results. It is shown in [136, Lemma 2.8] that the general position number of a maximal outerplane graph is bounded in terms of the maximum degree as: $\operatorname{gp}(G)\geq\lfloor\frac{2(\Delta(G)+1)}{3}\rfloor$. In the next result, maximal outerplane graphs with $\operatorname{gp}(G)=3$ are characterised.
@@
[136, Theorem 3.1] If $G$ is a maximal outerplane graph with $n(G)\geq 7$, then $\operatorname{gp}(G)=3$ if and only if $G$ is a straight linear $2$-tree.
In [136, Theorem 3.4], it was proved that for any maximal outerplane graph $G$ of order $n\geq 6$, $\operatorname{gp}(G)\leq\left\lfloor\frac{2n}{3}\right\rfloor$. The paper [136] then focuses on all maximal outerplane graphs that attain this bound.
@@ 2.3.3 Additional exact results
The general position numbers of glued binary trees $GT(r)$, $r\geq 2$, and of Beneš networks $BN(r)$, $r\geq 1$, were determined in [91] and [93] respectively. We do not give the definitions of these networks here, but recall that if $r\geq 2$, then $\operatorname{gp}(GT(r))=2^{r}$ [91, Proposition 3.8], and that if $r\geq 1$, then $\operatorname{gp}(BN(r))=2^{r+1}$ [93, Theorem 4.1]. Similar research has been done in [107, 108, 109]. In [107], the general position number is determined for hyper trees and shuffle hyper trees. In [108], the general position number is obtained for hexagonal networks, honeycomb networks, silicate networks, and oxide networks, while [109] reports the general position number of butterfly networks.
@@
[144, Theorem 2.5] If $G$ is a cactus graph with $k$ cycles and $t$ pendant edges, then $\operatorname{gp}(G)\leq\max\{3,2k+t\}$.
Moreover, [144, Theorem 2.5] also characterises the class of cactus graphs that attain equality in Theorem 2.18 as the cactus graphs in which all cycles are ‘good’; see the paper for the definition of a good cycle.
If $p\geq 3$ and $n\geq 1$, then the Sierpiński graph $S_{p}^{n}$ has vertex set $V(S_{p}^{n})=\{0,1,\ldots,p-1\}^{n}$, whilst vertices $i_{1}\cdots i_{n}$ and $j_{1}\cdots j_{n}$ are adjacent if there exists an index $h\in[n]$, such that (i) if $t<h$, then $i_{t}=j_{t}$, (ii) $i_{h}\neq j_{h}$, and (iii) if $t>h$, then $i_{t}=j_{h}$ and $j_{t}=i_{h}$, see [68]. In [118, Corollary 3.2] it is proved that if $p\geq 3$ then,
[…]
where $\operatorname{gp}(G)$ indicates the number of $\operatorname{gp}$-sets of a graph $G$. Moreover, [118, Corollary 4.5] asserts that if $n\geq 2$, then $\operatorname{gp}(S_{3}^{n})=3^{n-2}+3$.
@@ 2.4.1 Cartesian product
General position sets have been extensively researched so far on Cartesian products. As a first result, it was proved in [93, Corollary 3.2] that $\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4$, where $P_{\infty}$ is the two-way infinite path. In the same paper it was further proved that $10\leq\operatorname{gp}(P_{\infty}^{\operatorname{\,\square\,},3})\leq 16$, where $G^{\operatorname{\,\square\,},n}$ denotes the $n$-fold Cartesian product of $G$. These results were completed by the following result, where the Erdős-Szekeres Theorem on monotone sequences was applied to prove the upper bound.
@@
[78, Theorem 1] If $n\in{\mathbb{N}}$, then $\operatorname{gp}(P_{\infty}^{\operatorname{\,\square\,},n})=2^{2^{n-1}}$.
For general Cartesian products, we have the following lower bound.
@@
[56, Theorem 3.1] If $G$ and $H$ are connected graphs, then
@@
Tian and Xu followed with the following related result, where a graph is a generalised complete graph if it is obtained by the join of an isolated vertex with a disjoint union of one or more complete graphs.
Theorem 2.23.
[134, Theorems 3.7 and 3.8] If $G$ and $H$ are connected graphs with $n(G)\geq 3$ and $n(H)\geq 3$, then
$\operatorname{gp}(G\operatorname{\,\square\,}H)\leq n(G)+n(H)-2\,.$
Moreover, the equality holds if and only if both $G$ and $H$ are generalised complete graphs.
@@
[134, Theorem 3.5] If $G$ is a connected, bipartite graph with $n(G)\geq 3$, then $\operatorname{gp}(G\operatorname{\,\square\,}K_{2})=\operatorname{gp}(G)$ if and only if G is a complete bipartite graph.
@@
[56, Theorem 3.2] If $k\geq 2$ and $n_{1},\ldots,n_{k}\geq 2$, then
@@
    [137, Theorem 2.1] $\operatorname{gp}(T\operatorname{\,\square\,}T^{\prime})=\operatorname{gp}(T)+\operatorname{gp}(T^{\prime})$.
@@
    [134, Theorem 3.2] $\operatorname{gp}(T\operatorname{\,\square\,}K_{m})=\operatorname{gp}(T)+\operatorname{gp}(K_{m})-1$.
From Theorem 2.21 we can deduce that if $r\geq 3$ and $s\geq 3$, then $\operatorname{gp}(P_{r}\operatorname{\,\square\,}P_{s})=4$. This result has been expanded as follows, where $\operatorname{gp}(G)$ indicates the number of $\operatorname{gp}$-sets of a graph $G$.
@@
[76, Theorem 2.1] If $2\leq r\leq s$, then
$\operatorname{gp}(P_{r}\operatorname{\,\square\,}P_{s})=\left\{\begin{array}[]{ll}\vskip 5.69054pt6;&r=s=2\,,\\ \vskip 5.69054pt\displaystyle{\frac{s(s-1)(s-2)}{3}};&r=2,s\geq 3\,,\\ \displaystyle{\frac{rs(r-1)(r-2)(s-1)(s-2)(r(s-3)-s+7)}{144}};&r,s\geq 3\,.\end{array}\right$.
@@
[83, Theorem 3.4] If $r$ and $s$ are integers with $r\geq s\geq 4$, then
@@
A formula for the general position number of the Cartesian product of a complete multipartite graph with a path is also given in [134, Theorem 3.6].
The $n$-dimensional hypercube $Q_{n}$, $n\geq 1$, is defined to be $K_{2}^{\operatorname{\,\square\,},n}$, the Cartesian product of $n$ copies of $K_{2}$. To determine $\operatorname{gp}(Q_{n})$ is a notoriously difficult problem. The only exact values known so far are $\operatorname{gp}(Q_{1})=\operatorname{gp}(Q_{2})=2$, $\operatorname{gp}(Q_{3})=4$, $\operatorname{gp}(Q_{4})=5$, $\operatorname{gp}(Q_{5})=6$, $\operatorname{gp}(Q_{6})=8$, and $\operatorname{gp}(Q_{7})=9$, see [83, Table 1]. Using a probabilistic construction, Körner [82] obtained general position sets in $Q_{n}$ of size $\frac{1}{2}\frac{2^{n}}{\sqrt{3^{n}}}$. He also noted that the problem of finding $\operatorname{gp}(Q_{n})$ is equivalent to finding the largest size of a $(2,1)$-separating system in coding theory. Defining
@@
[76, Theorem 5.1] If $G$ is a graph, then
@@ 2.4.2 Strong product and direct product
[81, Theorem 4.2, Corollary 4.1] If $G$ and $H$ are connected graphs, then
@@
[81, Proposition 4.3] If $G$ is a connected graph and $n\geq 1$, then $\operatorname{gp}(G\boxtimes K_{n})=n\cdot\operatorname{gp}(G)$. Moreover, if $\operatorname{gp}(G)=\omega(G\operatorname{{}_{SR}})$, then $\operatorname{gp}(G\boxtimes K_{n})=\omega((G\boxtimes K_{n})\operatorname{{}_{SR}})$.
@@ 2.4.3 Generalised lexicographic product
[81, Theorem 5.1] Let $G$ be a graph with $V(G)=\{g_{1},\ldots,g_{n}\}$ and let $k_{i}$, $i\in[n]$, be positive integers. If $S$ is a gp-set of $G$ that induces a complete subgraph of $G\operatorname{{}_{SR}}$, and $\min\{k_{i}:\ g_{i}\in S\}\geq\max\{k_{i}:\ g_{i}\notin S\}$, then
@@ 2.4.4 Sierpiński product
[…]
@@
[132, Theorem 3.2] If $G$ is a connected graph of order at least two and $u\in V(G)$, then
@@
[132, Theorem 4.2] If $G$ and $H$ are two connected graphs of order at least two, then
@@
If $G=K_{2}$ and $H$ is a connected graph with $n(H)\geq 2$, then $\underline{{\rm gp}}{{}_{\rm S}}(K_{2},H)=2\xi^{-}(H)$ and ${\rm gp}{{}_{\rm S}}(K_{2},H)=2\xi(H)$, see [132, Theorem 4.4]. Finally, for Sierpiński products of complete graphs the following is true.
@@ 2.5.1 Vertex- and edge-deleted subgraphs
[42, Theorem 3.1] If $x$ is a vertex of a graph $G$, then $\operatorname{gp}(G-x)\leq 2\operatorname{gp}(G)$. Moreover, the bound is sharp.
The sharpness examples for Theorem 2.39 provided in [42] are such that GxG-x is not connected. Constructions are also given which demonstrate that $\operatorname{gp}(G-x)$ can be much larger than $\operatorname{gp}(G)$ even when GxG-x is connected.
On the other hand, $\operatorname{gp}(G-x)$ cannot be bounded from below by a function of $\operatorname{gp}(G)$. For instance, consider the fan graphs $F_{n}$, $n\geq 3$, that is, the join between $P_{n}$ and $K_{1}$. Then $\operatorname{gp}(F_{n})=\lceil\frac{2n}{3}\rceil$ and $\operatorname{gp}(F_{n}-x)=2$, where $x$ is the vertex of $F_{n}$ of degree $n$. On the positive side, if $x$ lies in some $\operatorname{gp}$-set, then $\operatorname{gp}(G)-1\leq\operatorname{gp}(G-x)$, see [42, Proposition 3.3]. For edge-deleted subgraphs we have:
@@ 2.5.2 Complementary prisms
[98, Theorem 3.1] If $G$ is a connected graph, then $\operatorname{gp}(G\overline{G})\leq n(G)+1$, and if $G$ is disconnected, then $\operatorname{gp}(G\overline{G})\leq n(G)$.
In the next result graphs are characterised for which the first upper bound of Theorem 2.41 is achieved. To state it, we need the following two concepts, where the second represents the original idea that led to the $d$-positions sets which will be considered in Section 3.1. A vertex $v$ of a connected graph $G$ is a central vertex of $G$ if ${\rm ecc}_{G}(v)=\operatorname{rad}(G)$. A set $S$ of vertices in a graph $G$ is a $3$-general position set if no three vertices from $S$ lie on a common shortest path of length at most three. By $N_{G}^{2}(v)$ we denote the set of vertices at distance two from $v$ in $G$.
@@
[98, Theorem 3.4] Let $G$ be a graph with $n(G)\geq 2$ and such that both $G$ and $\overline{G}$ are connected. Then $\operatorname{gp}(G\overline{G})=n(G)+1$ if and only if $\operatorname{rad}(G)=2$ and there exists a central vertex $v$ such that
@@
    $\overline{N_{G}(v)}$ is a $3$-general position set in $\overline{G}$ and $N_{G}^{2}(v)$ is a $3$-general position set in $G$, and
@@
[98, Theorem 4.1] If $G$ is a graph, then
@@
It follows from Theorem 2.43 that $\operatorname{gp}(G\overline{G})\geq\max\{\operatorname{gp}_{3}(G),\operatorname{gp}_{3}(\overline{G})\}$. In [98, Theorem 4.2] it was proved that among connected graphs the equality holds here if and only if $G$ is a complete multipartite graph. The paper [98] considers several additional classes of graphs $G$ and the corresponding values of $\operatorname{gp}(G\overline{G})$: bipartite graphs (in particular trees, hypercubes and Cartesian products of paths), split graphs and block graphs. For instance, if $T$ is a tree, then
[…]